OpenStudy (anonymous):

prove cosx/1+sinx=1-sinx/cosx

OpenStudy (anonymous):

cosx/(1-sinx) = (1+sinx)/cosx =cosx/(1-sinx) =(cosx(1+sinx))/((1-sinx)(1+sinx)) =(cosx+cosxsinx)/(1-sin²x) =(cosx+cosxsinx)/(cos²x) =(1+sinx)/cosx

OpenStudy (paxpolaris):

\[{\cos x \over1+\sin x}={1-\sin x \over \cos x}\]

OpenStudy (paxpolaris):

do you follow @meganlohry ?

OpenStudy (paxpolaris):

\[LHS\\={\cos x \over 1+ \sin x}\\={\cos x \over 1+ \sin x}\times {\left( 1-\sin x \right)\over\left( 1-\sin x \right)}\]

OpenStudy (paxpolaris):

\[={\cos x\left( 1-\sin x \right)\over 1-\sin^2x}\]

OpenStudy (paxpolaris):

\[={\cos x\left( 1-\sin x \right)\over \cos^2x} \\={1-\sin x \over \cos x}\\=RHS\]